the length of a rectangle is 4ft longer than its width. If the perimeter of the rectangle is 44ft, find its area.
Answer by nerdybill(7384)

the length of a rectangle is 4ft longer than its width. If the perimeter of the rectangle is 44ft, find its area.
.
Let w = width
then
w+4 = length
.
2(w + w+4) = 44
w + w+4 = 22
2w+4 = 22
2w = 18
w = 9 feet (width)
.
Length:
w+4 = 9+4 = 13 feet
.
Area:
9 * 13 = 117 square feet